OpenStudy (anonymous):

Write the equation in slope-intercept form of the line that has a slope of 3 and contains the point (2, 5).@Savannah_Noelle

OpenStudy (anonymous):

You first have to use the point-slope formula for the equation of a line: y - y1 = m(x - x1) where "m" is the slope and you use point (x1, y1) which is (2, 5) y - 5 = 3(x - 2) Now you can put it in slope-intercept form: y = mx + b y = 3x - 1

OpenStudy (anonymous):

Good luck to you in all of your studies and thx for the recognition! @ShoriteThug

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

uw!

OpenStudy (anonymous):

Write the equation in slope-intercept form of the line that has a slope of -3 and contains the point (4, -5).

OpenStudy (anonymous):

I'll do that one here since you wrote the question already, but if you have a 3rd question, please start a new thread. You first have to use the point-slope formula for the equation of a line: y - y1 = m(x - x1) where "m" is the slope and you use point (x1, y1) which is (4, -5) y - (-5) = -3(x - 4) Now you can put it in slope-intercept form: y = mx + b y = -3x + 7

OpenStudy (anonymous):

All good now, @ShoriteThug ?

OpenStudy (anonymous):

yes Write the equation in slope-intercept form of the line that has a slope of 5 and y-intercept of -3.

OpenStudy (anonymous):

I don't know if you missed it, but I asked you to write that in a new thread.

OpenStudy (anonymous):

i still dont get it

OpenStudy (anonymous):

It's ok if you still don't get it, but do write these problems in a new thread each time. Either I or another tutor can help you, but: new thread.